OpenStudy (anonymous):

how do i find the partial fraction decompisition -20x/3x^2+2x-8

OpenStudy (anonymous):

does the denominator factor?

OpenStudy (anonymous):

oh yes, it does \[-\frac{20x}{(3x-4)(x+2)}=\frac{a}{3x-4}+\frac{b}{x+2}\] the minus sign out front is kind of annoying, it would be easier to compute \[\frac{20x}{(3x-4)(x+2)}=\frac{a}{3x-4}+\frac{b}{x+2}\] then stick a minus sign outside the result

OpenStudy (anonymous):

what to do it the snap way?

OpenStudy (anonymous):

what do i do with the numerator?

OpenStudy (anonymous):

not sure what you mean your job is to find \(a\) and \(b\)

OpenStudy (anonymous):

oh i see, okay thank you

OpenStudy (anonymous):

do you know how to do it? there is a rather easy method for this one

OpenStudy (anonymous):

i would love to learn the easier method because i have no clue what to do

OpenStudy (anonymous):

ok lets do it the donkey way first, then we can do it the snap way. it amounts to the same thing

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{20x}{(3x-4)(x+2)}=\frac{a}{3x-4}+\frac{b}{x+2}\] is the start. now add up on the right pay no attention to the denominator because they will be the same you get \[20x=a(x+2)+b(3x-4)\] this has to be true not matter what \(x\) is , so pick \(x=-2\) to solve for \(b\) literally you get \[20\times -2=a(-2+2)+b(3\times -2-4)\] but you should really start with \[-40=b(-6-4)\] this tells you \(-40=-10b\implies b=10\)

OpenStudy (anonymous):

oops that was wrong \[-40=-10b\implies b=4\] sorry

OpenStudy (anonymous):

now repeat the process, this time with the value of \(x\) that will make \(3x-4=0\) namely pick \(x=\frac{4}{3}\)

OpenStudy (anonymous):

ok, im a little confused, how did you get rid of the (3x-4)(x+2) that was in the denomenator earlier, with the 20x in the numerator, to get 20x= a(3x-4)+b(x+2)

sam (.sam.):

\[\begin{array}{l} \text{factor the denominator into} \\ (x+2) (3 x-4) \\ \text{Then the partial fraction expansion is of the form} \\ -\frac{20 x}{(2+x) (-4+3 x)}=\frac{A}{x+2}+\frac{B}{3 x-4} \\ \text{Multiply both sides by }(x+2) (3 x-4)\text{ and simplify} \\ -20 x=A (3 x-4)+B (x+2) \\ \text{You need to make either A or B zero, so, try let x=-2,B will be zero,} \\ -20(-2)=A(3(-2)-4) \\ A=-4 \\ \text{Now do the same for A, let x=4/3, Then A will be zero, solve for B}\\\end{array}\]

OpenStudy (anonymous):

its the simplifying that im struggling with, it seems that everytime i simplify i end up with -20x/(3x-4)(x+2)=a(x+2)+b(3x-4)/3(x-4)(x+2)

sam (.sam.):

You just multiply both sides by (2+x)(-4+3x)

OpenStudy (anonymous):

okay thanks

sam (.sam.):

\[-\frac{20 x}{(2+x) (-4+3 x)}=\frac{A}{x+2}+\frac{B}{3 x-4}\] \[-\frac{20 x}{(2+x) (-4+3 x)}(2+x) (-4+3 x)=\frac{A}{x+2}(2+x) (-4+3 x)+\frac{B}{3 x-4}(2+x) (-4+3 x)\] \[-\frac{20 x}{\cancel{(2+x) (-4+3 x)}}\cancel{(2+x) (-4+3 x)}=\frac{A}{\cancel{x+2}}\cancel{(2+x)} (-4+3 x)+\frac{B}{\cancel{3 x-4}}(2+x) \cancel{(-4+3 x)}\] \[-20x=A(-4+3x)+B(2+x)\]

OpenStudy (anonymous):

thank you soooo much , i get it now.